OpenStudy (anonymous):

Use a triple integral to find the volume of the given solid: The solid enclosed by the cylinder x^2 + y^2 = 9 and the planes y + z = 5 and z=1. Please explain step by step.

OpenStudy (anonymous):

T_T. Calc 3. brutal man. use circular coordinates i think? turn x^2+y^2 = 9 into r^2 = 9. r will be from 0 to 3. Thats one bound. z = 5-y. another bound is z goes from 1 to (5-y). thats a start, Im having a hard time remembering the variables you use for this type.

OpenStudy (anonymous):

the last variable u use is theta i belive. from 0 to 2pi. turn y into rsin(theta). The three will be 0->3, 0->2pi, 1->(5-y) S(r*dz*d(theta)*dr). Something similiar. This will not work out...at least I dont think it will. sorry

OpenStudy (anonymous):

first sketch to have a better view on the problem so we'll have:   as we can see the upper and lower limits for our integral will be \[-3\le x \le3\] \[-\sqrt{9-x^2} \le y \le \sqrt{9-x^2}\] \[1 \le z \le 5-y\] so our triple integral will become: \[\int\limits_{-3}^{3}\int\limits_{-\sqrt{9-x^2}}^{\sqrt{9-x^2}}\int\limits_{1}^{5-y}dV\]